4. Sally invested a sum of money at 10% interest, compounded monthly. She let the sum accumulate interest for 12 years and now has $15,000. What was Sally's initial amount that she invested?
Answer #1

Answer:

Given

Final value of Investment F=$15,000

Interest Rate=10% per annum

r=10%/12=0.833% per month

Years =12 years

Periods n=12*12=144 months

Let Sally invested X amount.

So F=X*(1+r)^n

X=F/(1+r)^n

X=15000/(1+0.833%)^144

X=$4542.60